Suppose we are told that the speed, v , of a rotating particle depends on the radius, r , and the frequency f , in the form v=2πr^a / f^b . If this is to be a dimensionally homogeneous equation, the values of a and b are...


1
Expert's answer
2020-11-02T09:24:22-0500

The speed has the dimensions "m\/s", the radius - "m", and the frequency - "Hz = 1\/s". Thus, have the following equation:


"\dfrac{m^a}{(1/s)^b} = \dfrac{m}{s}"

Comparing the powers on the both sides of the equation, one can derive the values of "a" and "b":


"a =1\\
b = -1"

Answer. a = 1, b = -1.